Just checking to see if I did this problem correctly.
Find the equation of the line that goes through the points (-4,1) and (-1,4)
I found the slope of -1/-1 or -1
y = x + b
4 = -1 + b
5 = b
answer y = x + 5 ??

Found 2 solutions by Alan3354, nerdybill:
Answer by Alan3354(69443)

Just checking to see if I did this problem correctly.
Find the equation of the line that goes through the points (-4,1) and (-1,4)
============
I found the slope of -1/-1 or -1 ************* -1/-1 = +1
------------
y = mx + b
4 = 1*-1 + b
b = 5
--------
y = x + b
4 = -1 + b
5 = b
answer y = x + 5 Is correct.
-------------
You can check your answer by subbing the points.

Answer by nerdybill(7384)

Your answer is correct. However, I'm not sure I followed how you got there.
.
slope of (-4,1) and (-1,4):
m = (y2-y1)/(x2-x1)
m = (4-1)/(-1-(-4))
m = (4-1)/(-1+4)
m = 3/3
m = 1
.
plug slope (above) and one point (-4,1) into "point-slope" form:
y - y1 = m(x - x1)
y - 1 = 1(x - (-4))
y - 1 = 1(x + 4)
y - 1 = x + 4
y = x + 5